# Artificial Intelligence for Microbiology and Microbiome Research

### v1

@@ Abstract
Advancements in artificial intelligence (AI) have transformed many scientific fields, with microbiology and microbiome research now experiencing significant breakthroughs through machine learning and deep learning applications. This review provides a comprehensive overview of AI-driven approaches tailored for microbiology and microbiome studies, emphasizing both technical advancements and biological insights. We begin with an introduction to foundational AI techniques, including primary machine learning paradigms and various deep learning architectures, and offer guidance on choosing between machine learning and deep learning methods based on specific research goals. The primary section on application scenarios spans diverse research areas, from taxonomic profiling, functional annotation & prediction, microbe-X interactions, microbial ecology, metabolic modeling, precision nutrition, clinical microbiology, to prevention & therapeutics. Finally, we discuss challenges unique to this field, including the balance between interpretability and complexity, the "small n, large p" problem, and the critical need for standardized benchmarking datasets to validate and compare models. Together, this review underscores AI’s transformative role in microbiology and microbiome research, paving the way for innovative methodologies and applications that enhance our understanding of microbial life and its impact on our planet and our health.
@@ 1 Introduction
It is not a big surprise that disrupted microbial communities (or microbiomes) can have a huge impact on our planet and ourselves. Indeed, agricultural practices, such as excessive use of chemical fertilizers and pesticides, can disrupt soil microbiomes, leading to reduced soil fertility and increased vulnerability to erosion [6]. Runoff containing pollutants and antibiotics can significantly disrupt the microbiomes of freshwater and marine ecosystems, leading to changes in water quality and impacting the health of aquatic life by altering the natural balance of microbial communities within the environment; this can potentially promote the growth of harmful bacteria and disrupt critical ecological processes like nutrient cycling [7, 8]. Many human diseases have been associated with disrupted microbiomes, including acne, eczema, dental caries, obesity, malnutrition, inflammatory bowel disease, asthma/allergies, hardening of arteries, colorectal cancer, type 2 diabetes, as well as neurological conditions such as autism, anxiety, depression, and post-traumatic stress disorder, etc [9, 10]. Gaining a deeper understanding of the activities of microbial communities, both within and around us, can greatly benefit our health and the health of our planet. This explains why in the past decades the microbiome has been a very active research topic in microbiology.
@@
Artificial Intelligence (AI) focuses on creating intelligent machines that can execute tasks that usually need human intelligence. AI emerged as an academic discipline at the 1956 Dartmouth conference, shaped by pioneering work by Warren McCulloch, Walter Pitts, and Alan Turing on neural networks and machine intelligence. At first, AI research concentrated on symbolic reasoning, including early applications in biomedicine, such as the MYCIN expert system for diagnosing bacterial infections. Meanwhile, machine learning developed, showcasing algorithms that improved through data training. Despite early excitement and positive forecasts, the pace of AI advancement decelerated over the following decades, hindered by hardware constraints and unmet expectations, leading to a period known as "AI winter." However, the domain continued to progress, incorporating probabilistic methods to manage uncertainty. In around 2010, a new phase in AI emerged, fueled by breakthroughs in deep learning frameworks, the advent of powerful hardware (e.g., GPUs), open-source software tools, and greater access to extensive datasets (e.g., ImageNet [11]). In 2012, significant breakthroughs occurred when AlexNet (a deep learning architecture based on the convolutional neural network) surpassed preceding machine learning methodologies in visual recognition [12]. The subsequent innovations, particularly the Transformer (a deep learning architecture initially developed for machine translation) introduced in 2017 [13], triggered an "AI boom" marked by considerable investment. This surge in investment led to a wide range of AI applications by the 2020s, accompanied by increasing concerns regarding its societal implications and the pressing need for regulatory measures.
@@
In this article, we review the application of various AI techniques in microbiology and microbiome research. We will focus on the applications of machine learning, particularly deep learning techniques. Traditional microbiologists excel in image analysis skills for identifying pathogens in Gram stains, ova and parasite preparations, blood smears, and histopathologic slides. They classify colony growth on agar plates for assessment. AI advances in computer vision can automate these processes, supporting timely and accurate diagnoses [14, 15]. Advances in sequencing technologies, especially next-generation sequencing, enable substantial numbers of samples to be processed rapidly and cost-efficiently [16]. The accessibility of large-scale microbiome datasets propelled the development of numerous AI (especially machine learning or deep learning) approaches in microbiome studies, as reviewed previously [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 29, 43, 44, 45, 46, 47, 48, 49, 50, 29, 51]. However, a comprehensive review of existing applications of AI techniques in microbiology and microbiome research is still lacking. This review article aims to fill this gap. The following sections are organized as follows. We first briefly describe various AI subfields, focusing on machine learning and the three basic machine learning paradigms. Next, we elaborate on the different deep learning techniques categorized under the three primary machine learning paradigms. Then, we systematically review the various applications of AI techniques in microbiology and microbiome research. Finally, we will present an outlook on the future directions of AI for microbiology and microbiome research.
@@ 2.1 Learning Paradigms
Supervised learning involves using labeled datasets, where each data point is linked to a class label. The algorithms in this approach aim to create a mathematical function that connects input features to the expected output values, relying on these labeled instances. Common uses include classification and regression. Classical machine learning methods for classification/regression include Logistic Regression, Naïve Bayes, Support Vector Machine (SVM), Random Forest, Extreme Gradient Boosting (XGBoost), etc. Those methods have been heavily used in microbiology and microbiome research.
@@
In unsupervised learning, algorithms analyze unlabeled data to detect patterns and relationships without any defined categories. This process uncovers similarities in the dataset and includes techniques like clustering, dimensionality reduction, and association rules mining. Classical unsupervised learning methods include k-means clustering, Principal Component Analysis (PCA), Principal Coordinate Analysis (PCoA), and t-distributed stochastic neighbor embedding (t-SNE) for dimension reduction, and the Apriori algorithm for association rules mining. Among them, PCoA is a commonly used tool in microbiome data analysis, particularly valuable for visualizing and interpreting the differences in microbial community composition between samples.
@@
Reinforcement learning focuses on enabling intelligent agents to learn through trial-and-error in a dynamic environment to maximize their cumulative rewards [54, 55, 56]. Without labeled datasets, these agents make decisions to maximize rewards, engaging in autonomous exploration and knowledge acquisition, which is crucial for tasks that are difficult to program explicitly.
@@
Integrating these paradigms can often lead to better outcomes. For instance, semi-supervised learning finds a middle ground by utilizing a small set of labeled data alongside a larger collection of unlabeled data. This method harnesses the strengths of both supervised and unsupervised learning, making it a cost-effective and efficient way to train models when labeled data is scarce. In situations where obtaining high-quality labeled data is difficult, self-supervised learning presents a viable alternative [57]. In this framework, models are pre-trained on unlabeled data, with labels generated automatically in subsequent iterations. Self-supervised learning effectively converts unsupervised machine learning challenges into supervised tasks, improving learning efficiency.
@@
Transfer learning is another interesting machine learning technique, which involves taking a pre-trained model on a large dataset and fine-tuning it on a smaller, task-specific dataset [58, 59]. This approach leverages the knowledge acquired by the model during pre-training to improve performance on a new task. Transfer learning can be applied within both supervised and unsupervised learning paradigms, meaning it can utilize knowledge learned from either labeled or unlabeled data depending on the situation; essentially, transfer learning "transfers" the learned representations from one task to another, regardless of whether the original task was supervised or unsupervised.
@@
Note that both self-supervised learning and transfer learning leverage pre-trained models to improve performance on new tasks, but the key difference is that self-supervised learning generates its own labels, often called “pseudo-labels”, from unlabeled data during the pre-training phase, while transfer learning relies on existing labeled or unlabeled data for pre-training. Both self-supervised learning and transfer learning are extensively used in the training of large language models (LLMs), with self-supervised learning often being the primary method for pre-training on massive amounts of unlabeled data, while transfer learning allows the pre-trained model to be adapted to specific downstream tasks with fine-tuning on smaller labeled datasets. LLMs tailored for biology, e.g., genomic and protein language models [60, 61, 62, 63, 64], have numerous applications in microbiology and microbiome research. These models, trained on vast amounts of biological sequence data, can generate insights and predictions that are valuable across various areas in microbiology and microbiome research, as we discuss later.
2.2 Deep learning techniques
@@
Figure 1: A taxonomy of deep learning techniques. Figure adapted from Ref [70]. MLP: Multi-Layer Perceptron; CNN: Convolutional Neural Network; ResNet: Residual Neural Network; GCN: Graph Convolutional Network; GAT: Graph Attention Network; RNN: Recurrent Neural Network; LSTM: Long Short-Term Memory; GRU: Gated Recurrent Unit; SAT: Structure-Aware Transformer; GAN: Generative Adversarial Network; AE: Auto-Encoder; SAE: Sparse Autoencoder; DAE: Denoising Autoencoder; CAE: Contractive Autoencoder; VAE: Variational Autoencoder; SOM: Self-Organizing Map; RBM: Restricted Boltzmann Machine; DBN: Deep Belief Network; DRL: Deep Reinforcement Learning.
Based on the three primary machine learning paradigms, deep learning can be broadly divided into three major categories (Fig.1). The first category includes deep networks for supervised or discriminative learning, such as Multi-Layer Perceptron (MLP), Convolutional Neural Network (CNN) and their variants, Recurrent Neural Network (RNN) and their variants, as well as the Transformer. Roughly speaking, RNN propagates information through all hidden states in a sequential way, while CNN takes local information in developing each representation. By contrast, Transformer develops global contextual embedding via self-attention [13], which enables models to dynamically determine the relative importance of various words in a sequence, improving the ability to capture long-range dependencies. Another big advantage of Transformer is its easy parallelism. Unlike RNN, the Transformer can process entire sequences in parallel, which allows us to use GPUs for training. This significantly reduces the training time, and allows the use of very large models, often with hundreds of billions of parameters. These two advantages explain why the Transformer has facilitated so many LLMs, e.g., BERT, T5, GPT, PaLM, Gemini, and has revolutionized AI. As we will see later, all those deep network architectures in the first category (i.e., MLP, CNN, RNN, and Transformer), which were originally used for supervised learning, have been widely used in microbiome research.
The second category includes deep networks for unsupervised or generative learning, such as Generative Adversarial Network (GAN), Autoencoder (AE) and its variants, Self-Organizing Map (SOM), Restricted Boltzmann Machine (RBM), and Deep Belief Network (DBN). GAN is a very popular neural network architecture in recent years [71]. This architecture uses the idea of game theory to train two neural networks to compete with each other, thereby generating more realistic new data from a given training data set. AE is also a very common unsupervised neural network model, which can learn the latent features of the input data (called encoding), and at the same time use the learned features to reconstruct the original input data (called decoding) [72]. There are many variants of AE. Among them, the variational autoencoder (VAE) is probably the most famous one. VAE uses a probabilistic framework. Instead of mapping the input to a single point in the latent space, VAE maps the input to a distribution on the latent space, allowing for more flexible and expressive data representation [73]. As we will see later, both GAN and AE have been widely used in microbiome research. The other three models (SOM, RBM, and DBN) have not.
The third category includes deep networks for hybrid learning and relevant other tasks. There are three kinds of hybrid learning models: (1) An integration of different generative (or discriminative) models to extract more meaningful and robust features, e.g., CNN+LSTM, AE+GAN; (2) An integration of a generative model followed by a discriminative model, e.g., DBN+MLP, GAN+CNN, AE+CNN, etc; (3) An integration of generative or discriminative model followed by a non-deep learning classifier, e.g., AE+SVM, CNN+Random Forest, etc. As we will see later, all three hybrid learning models have been widely used in microbiome research. This category also includes Deep Reinforcement Learning (DRL). DRL is a subfield of machine learning that combines reinforcement learning and deep learning. Reinforcement Learning helps agents learn decision-making through trial and error. DRL improves this by using deep learning to extract decisions from unstructured data without manual state space engineering. DRL algorithms can take in very large inputs (e.g., an image of the raw board state and the history of states) and decide what actions to perform to optimize an objective (e.g., winning the game). A famous DRL algorithm is AlphaGo Zero, learning from playing the ancient Chinese game of Go without using any human knowledge [74]. So far, applications of DRL techniques in microbiome research are still very rare.
2.3 When to Use Machine learning vs. Deep learning?
@@
We do not always need fancy deep learning techniques for microbiology and microbiome research. Sometimes we do not need deep learning at all. Logistic Regression or Random Forest might work very well. Choosing between deep learning and traditional machine learning methods depends on data characteristics, the specific problem at hand, available computational resources, and the need for model interpretability. Traditional methods are generally preferred for smaller, structured datasets and scenarios requiring interpretability (such as clinical applications), while deep learning excels with large, unstructured datasets and complex tasks requiring high performance.
@@
If we decide to apply or develop deep learning methods to solve our problem, there is a general procedure [75]. First, we need to choose the appropriate performance metrics (e.g., Accuracy, Precision, Recall, F1-score, AUROC, AUPRC). Second, we need to find the default baseline deep learning models based on the data structure. For supervised learning tasks that involve fixed-size vector inputs, it is advisable to utilize a feedforward network featuring fully connected layers (e.g., MLP). If the input possesses a known topological structure, such as images or graphs, opting for CNN or its variants (e.g., graph convolutional network (GCN)) is recommended. When dealing with inputs or outputs that form sequences, we should consider using RNN and its variant (e.g., LSTM or GRU) or Transformer. 1D CNN or temporal convolutional network (TCN) might also work. Depending on the task, a hybrid deep learning model could also be considered. Third, we need to establish a reasonable end-to-end system, which involves choosing the appropriate optimization algorithm (e.g., SGD with momentum, Adam) and incorporating regularization (via early stop, dropout, or batch normalization). Finally, we need to measure the performance and determine how to improve it. We can either gather more training data or tune hyperparameters (e.g., learning rate, number of hidden units) via grid search or random search.

### v2

@@ Abstract
Advancements in artificial intelligence (AI) have transformed many scientific fields, with microbiology and microbiome research now experiencing significant breakthroughs through machine learning applications. This review provides a comprehensive overview of AI-driven approaches tailored for microbiology and microbiome studies, emphasizing both technical advancements and biological insights. We begin with an introduction to foundational AI techniques, including primary machine learning paradigms and various deep learning architectures, and offer guidance on choosing between traditional machine learning and sophisticated deep learning methods based on specific research goals. The primary section on application scenarios spans diverse research areas, from taxonomic profiling, functional annotation & prediction, microbe-X interactions, microbial ecology, metabolic modeling, precision nutrition, clinical microbiology, to prevention & therapeutics. Finally, we discuss challenges in this field and highlight some recent breakthroughs. Together, this review underscores AI’s transformative role in microbiology and microbiome research, paving the way for innovative methodologies and applications that enhance our understanding of microbial life and its impact on our planet and our health.
@@ 1 Introduction
It is not a big surprise that disrupted microbial communities (or microbiomes) can have a huge impact on our planet and ourselves. Indeed, agricultural practices, such as excessive use of chemical fertilizers and pesticides, can disrupt soil microbiomes, leading to reduced soil fertility and increased vulnerability to erosion [6]. Runoff containing pollutants and antibiotics can significantly disrupt the microbiomes of freshwater and marine ecosystems, leading to changes in water quality and impacting the health of aquatic life by altering the natural balance of microbial communities within the environment; this can potentially promote the growth of harmful bacteria and disrupt critical ecological processes like nutrient cycling [7, 8]. Many human diseases have been associated with disrupted microbiomes, including acne, eczema, dental caries, obesity, malnutrition, inflammatory bowel disease, asthma/allergies, hardening of arteries, colorectal cancer, type 2 diabetes, as well as neurological conditions such as autism, anxiety, depression, and post-traumatic stress disorder, etc [9, 10, 11]. Gaining a deeper understanding of the activities of microbial communities, both within and around us, can greatly benefit our health and the health of our planet. This explains why in the past decades the microbiome has been a very active research topic in microbiology.
@@
Artificial Intelligence (AI) focuses on creating intelligent machines that can execute tasks that usually need human intelligence. AI emerged as an academic discipline at the 1956 Dartmouth conference, shaped by pioneering work by Warren McCulloch, Walter Pitts, and Alan Turing on neural networks and machine intelligence. At first, AI research concentrated on symbolic reasoning, including early applications in biomedicine, such as the MYCIN expert system for diagnosing bacterial infections. Meanwhile, machine learning developed, showcasing algorithms that improved through data training. Despite early excitement and positive forecasts, the pace of AI advancement decelerated over the following decades, hindered by hardware constraints and unmet expectations, leading to a period known as “AI winter." However, the domain continued to progress, incorporating probabilistic methods to manage uncertainty. In around 2010, a new phase in AI emerged, fueled by breakthroughs in deep learning frameworks, the advent of powerful hardware (e.g., GPUs), open-source software tools, and greater access to extensive datasets (e.g., ImageNet [12]). In 2012, significant breakthroughs occurred when AlexNet (a deep learning architecture based on the convolutional neural network) surpassed preceding machine learning methodologies in visual recognition [13]. The subsequent innovations, particularly the Transformer (a deep learning architecture initially developed for machine translation) introduced in 2017 [14], triggered an “AI boom" marked by considerable investment. This surge in investment led to a wide range of AI applications by the 2020s, accompanied by increasing concerns regarding its societal implications and the pressing need for regulatory measures.
@@
Traditional microbiologists excel in image analysis skills for identifying pathogens in Gram stains, ova and parasite preparations, blood smears, and histopathologic slides. They also characterize colony growth on agar plates to identify microorganisms and assess their properties, such as antibiotic susceptibility or ecological roles, often complementing these techniques with modern molecular methods. AI advances in computer vision can automate these processes, supporting timely and accurate diagnoses [15, 16]. Advances in sequencing technologies, especially next-generation sequencing, enable substantial numbers of samples to be processed rapidly and cost-efficiently [17]. The accessibility of large-scale microbiome datasets propelled the development of numerous AI (especially machine learning) approaches in microbiome studies [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 29, 43, 44, 45, 46, 47, 48, 49, 50, 29, 51]. However, a comprehensive review of existing applications of AI techniques in microbiology and microbiome research is still lacking. This review article aims to fill this gap. The following sections are organized as follows. We first briefly describe various AI subfields, focusing on machine learning and the three basic machine learning paradigms. Next, we elaborate on the different deep learning techniques categorized according to their underlying inductive biases and learning objectives. Then, we systematically review the various applications of AI techniques in microbiology and microbiome research. Finally, we discuss outstanding challenges and future directions of AI for microbiology and microbiome research.
@@ 2.1 Machine Learning Paradigms
Supervised learning involves using labeled datasets, where each data point is linked to a class label. The algorithms in this approach aim to create a mathematical function that connects input features to the expected output values, relying on these labeled instances. Common uses include classification and regression. Traditional machine learning methods for classification/regression include Logistic Regression, Naïve Bayes, Support Vector Machine (SVM), Random Forest (RF), Extreme Gradient Boosting (XGBoost), etc. Those methods have been heavily used in microbiology and microbiome research.
@@
In unsupervised learning, algorithms analyze unlabeled data to detect patterns and relationships without any defined categories. This process uncovers similarities in the dataset and includes techniques like clustering, dimensionality reduction, and association rules mining. Classical unsupervised learning methods include k-means clustering, Principal Component Analysis (PCA), Principal Coordinate Analysis (PCoA), t-distributed stochastic neighbor embedding (t-SNE), and Uniform Manifold Approximation and Projection (UMAP) for dimension reduction, and the Apriori algorithm for association rules mining. Among them, PCoA is a commonly used tool in microbiome data analysis, particularly valuable for visualizing and interpreting the differences in compositions of different samples.
@@
Reinforcement learning (RL) focuses on enabling intelligent agents to learn through trial-and-error in a dynamic environment to maximize their cumulative rewards [54]. Without labeled datasets, these agents make decisions to maximize rewards, engaging in autonomous exploration and knowledge acquisition, which is crucial for tasks that are difficult to program explicitly. RL has seen limited adoption in microbiology and microbiome research to date, despite its proven efficacy in other domains like robotics, game playing, and drug optimization. This scarcity stems from challenges inherent to microbial systems—such as high-dimensional data from metagenomics, complex community dynamics, and the need for interpretable models in biological contexts—making RL’s trial-and-error paradigm harder to adapt than supervised learning approaches like classification or clustering, which dominate current microbiome studies.
@@
Integrating these paradigms can often lead to better outcomes. For instance, semi-super-vised learning finds a middle ground by utilizing a small set of labeled data alongside a larger collection of unlabeled data. This method harnesses the strengths of both supervised and unsupervised learning, making it a cost-effective and efficient way to train models when labeled data is scarce. In situations where obtaining high-quality labeled data is difficult, self-supervised learning presents a viable alternative [55]. In this framework, models are pre-trained on unlabeled data, with labels generated automatically in subsequent iterations. Self-supervised learning effectively converts unsupervised machine learning challenges into supervised tasks, improving learning efficiency.
@@
Transfer learning is another powerful machine learning technique, which involves taking a pre-trained model on a large dataset and fine-tuning it on a smaller, task-specific dataset [56, 57]. This approach leverages the knowledge acquired by the model during pre-training to improve performance on a new task. Transfer learning can be applied within both supervised and unsupervised learning paradigms, meaning it can utilize knowledge learned from either labeled or unlabeled data depending on the situation; essentially, transfer learning "transfers" the learned representations from one task to another, regardless of whether the original task was supervised or unsupervised.
@@
Note that both self-supervised learning and transfer learning leverage pre-trained models to improve performance on new tasks, but the key difference is that self-supervised learning generates its own labels, often called “pseudo-labels”, from unlabeled data during the pre-training phase, while transfer learning relies on existing labeled or unlabeled data for pre-training. Both self-supervised learning and transfer learning are extensively used in the training of large language models (LLMs), with self-supervised learning often being the primary method for pre-training on massive amounts of unlabeled data, while transfer learning allows the pre-trained model to be adapted to specific downstream tasks with fine-tuning on smaller labeled datasets. LLMs tailored for biology, e.g., genomic and protein language models [58, 59, 60, 61, 62, 63, 64, 65], have numerous applications in microbiology and microbiome research. These models, trained on vast amounts of biological sequence data, can generate insights and predictions that are valuable across various areas in microbiology and microbiome research.
2.2 Deep learning architectures
@@
Deep learning architectures can be broadly categorized according to their underlying inductive biases—the structural assumptions that guide how they process data. Models designed for tabular and dense features often rely on multilayer perceptrons (MLPs) and increasingly incorporate specialized transformer-based architectures such as TabTransformer [71], TabNet [72], and FT-Transformer [73] to better capture feature interactions. Convolutional neural networks (CNNs), e.g., LeNet-5 [74], AlexNet [13], ResNet [75], U-Net [76], ConvNeXt [77], introduce locality and translation-equivariance priors, making them particularly effective for images and other spatially structured inputs. Attention-based and Transformer architectures, e.g., Transformer [14], GPT [78], BERT [79], ViT [80], Perceiver [81], Llama [82], emphasize flexible, content-dependent interactions and now dominate natural language processing and are increasingly applied to vision and multimodal learning. For time-series and sequential data, recurrent neural networks (RNNs) [83, 84] and its variants: Long short-term memory (LSTM) [85] and Gated Recurrent Unit (GRU) [86] have historically been used. Receptance Weighted Key Value (RWKV) [87] is a hybrid model architecture that combines the efficient parallelizable training of Transformers with the efficient inference of RNNs. Originated in control systems engineering, state-space models (SSM), e.g., High-order Polynomial Projection Operators (HiPPO) [88], structured state space sequence (S4) model [89] and Mamba [90], offer efficient long-range sequence modeling through continuous-time and linear-state dynamics. They have proven to match the performance capacity of transformer-based models while offering superior speed and efficiency. Graph neural networks (GNNs) and its variants e.g., graph convolutional network (GCN) [91], graph attention network (GAT) [92], graph recurrent network (GRN) [93], incorporate relational inductive biases tailored to graph-structured inputs such as molecular graphs. By effectively encoding the structural information of a graph into the standard Transformer architecture, Graphormer [94] has attained excellent results on a broad range of graph representation learning tasks. Finally, neural operator models, e.g., Deep Operator Networks (DeepONets) [95], Fourier Neural Operator (FNO) [96], Laplace Neural Operator [97], extend deep learning to learn solution operators for differential equations, enabling data-driven modeling across scientific and engineering domains. Together, these architectural families illustrate how inductive biases shape modern deep learning systems to match the structural properties of diverse data modalities.
Deep learning models are fundamentally distinguished by their learning objectives, which define the mathematical target optimized during training. Discriminative objectives model the conditional probability $p(y|x)$ to perform classification or regression, exemplified by ResNet [75] for image classification and BERT [79] (with a classification head) for text categorization. Generative objectives learn the data distribution $p(x)$ or $p(x,y)$, split into explicit density models (e.g., Variational Autoencoders (VAEs) [98], Pixel Recurrent Neural Networks (PixelRNN) [99], and Normalizing Flows (NF) [100]) which provide tractable likelihoods) and implicit sampling models (e.g., Generative Adversarial Networks (GANs) [101] and Diffusion Models (including Denoising Diffusion Probabilistic Models (DDPMs) [102], Stable Diffusion [103] and Sora [104]) which generate high-fidelity samples without explicit densities. Contrastive/Self-Supervised objectives learn invariant representations by maximizing similarity between augmented views of the same instance, as in SimCLR [105], BYOL [106], and DINO [107] for images or SimCSE [108] for text. Reconstructive objectives minimize input-output divergence, typically via autoencoding, with Masked Autoencoders (MAE) and standard Autoencoders (AE) learning compressed latent representations. Predictive objectives forecast future tokens or frames in a sequence, powering large language models like GPT and LLaMA (next-token prediction) and video generation models like Phenaki [109]. These objectives are largely orthogonal to architecture and supervision, enabling flexible combinations in modern foundation models.
2.3 Traditional Machine Learning vs. Deep Learning: When to Use Each
@@
As we will see later, many deep learning architectures (e.g., MLP, CNN, RNN, Transformer, GAN, AE), their variants (e.g., LSTM, VAE, GCN, BERT), their hybrids (e.g., CNN+LSTM, AE+GAN; GAN+CNN, AE+CNN), and their hybrids with traditional machine learning models (e.g., AE+SVM, CNN+Random Forest) have been widely used in microbiology and microbiome research. However, we have to emphasize that we do not always need advanced deep learning techniques for microbiology and microbiome research. Sometimes we do not need deep learning at all. Logistic Regression or Random Forest might work very well. Choosing between deep learning and traditional machine learning methods depends on data characteristics, the specific problem at hand, available computational resources, and the need for model interpretability. Traditional methods are generally preferred for smaller, structured datasets and scenarios requiring interpretability (such as clinical applications), while deep learning excels with large, unstructured datasets and complex tasks requiring high performance.
@@
If we decide to apply or develop deep learning methods to solve our problem, there is a systematic approach [110]. First, we need to choose performance metrics tailored to the task and dataset characteristics, such as Silhouette Index, Adjusted Rand Index, or Normalized Mutual Information for clustering; FID or Inception Score for generative models; accuracy, precision, recall, F1-score, AUC-ROC, or AUC-PR for classification; MSE, MAE, RMSE, or R² for regression; and MASE or sMAPE for time series forecasting, ensuring alignment with the problem’s objectives and constraints. Second, we need to find the default baseline deep learning models based on the data structure. For supervised learning tasks that involve fixed-size vector inputs, it is advisable to utilize a feedforward network featuring fully connected layers (e.g., MLP) or transformer-based architectures (e.g., TabTransformer, TabNet, or FT-Transformer). If the input possesses a known topological structure, such as images or graphs, opting for CNN or its variants (e.g.,GCN) is recommended. CNNs excel in processing Euclidean data, like microbial imaging or spatial microbiome distributions, by leveraging local receptive fields and weight sharing, whereas GCNs are suited for non-Euclidean data, such as microbiome interaction graphs or phylogenetic trees, by aggregating relational information across nodes to model microbial community dynamics or taxonomic dependencies. When dealing with inputs or outputs that form sequences, we should consider using RNN and its variant (e.g., LSTM or GRU) or Transformer. 1D CNN or temporal convolutional network (TCN) might also work. Depending on the task, a hybrid deep learning model could also be considered. Third, we need to establish a reasonable end-to-end system, which involves choosing the appropriate optimization algorithm (e.g., SGD with momentum, Adam) and incorporating regularization (via early stop, dropout, or batch normalization). Finally, we need to measure the performance and determine how to improve it. We can either gather more training data or tune hyperparameters (e.g., learning rate, number of hidden units) via grid search or random search.
2.4 Data Types in AI Applications for Microbiome Research
At the intersection of AI and microbiology/microbiome research, the choice of data type is critical for uncovering microbial functions, interactions, and ecological dynamics. AI models—from classical machine learning to deep learning—leverage diverse omics layers to predict functions, classify taxa, and simulate communities. Core data types include metagenomics (DNA sequences of communities), metatranscriptomics (RNA expression), metaproteomics (protein profiles), metabolomics (microbial metabolites), and meta-epigenomics (epigenetic modifications). Together, these offer a multidimensional view that enables AI to detect complex relationships often missed by traditional analyses.
Equally important are the input formats. Raw sequencing data (e.g., FASTQ) are information-rich but computationally demanding. To reduce dimensionality, they can be transformed into k-mers or relative abundances from profiling tools, which integrate more easily with standard models but risk losing biological nuance. More advanced approaches use LLM embeddings (e.g., ESM-2 [111], ProtGPT2 [112], ProtBERT [62]), which encode sequences into dense vectors capturing higher-order relationships and enabling transfer learning.
The choice of data type and format shapes model performance and generalizability. Metagenomics offers richness but risks overfitting without normalization or augmentation. Metaproteomics and metabolomics add functional depth but suffer from missing values and batch effects unless carefully preprocessed. Raw sequences require heavy computation and advanced architectures, while relative abundances trade scalability for detail. LLM embeddings improve transferability across datasets but demand careful fine-tuning to avoid bias. Ultimately, hybrid, multi-omics integration—when harmonized properly—yields more robust and generalizable AI models, though it requires strong interdisciplinary expertise.